OpenStudy (love.kat123):

How many solutions will the following system of linear equations have? –2x + 4y = –7 y = -1/2x + 5 A. 0 B. 1 C. 2 D. infinite For which system of equations is (5, 3) the solution? A. 3x – 2y = 9 3x + 2y = 14 B. x – y = –2 4x – 3y = 11 C. –2x – y = –13 x + 2y = –11 D. 2x – y = 7 2x + 7y = 31 Solve the system of equations. 8x – 3y = 1 –2x + 3y = 11 A. (–1, –3) B. (–1, 3) C. (2, 5) D. (5, 2)

OpenStudy (love.kat123):

@Adjax @texaschic101 @anonymous_user

OpenStudy (anonymous):

for the 1st question only one solution will be there

OpenStudy (anonymous):

2nd question D is the answer reason substitute the values its simple or solve it by normal equations

OpenStudy (texaschic101):

1. you have to compare the slopes and the y intercepts. In y = mx + b form, the slope is in the m position and the y intercept is in the b position. -2x + 4y = -7 4y = 2x - 7 y = 1/2x - 7/4 slope = 1/2 and y int = -7/4 y = -1/2x + 5 slope = -1/2 and y int = 5 if slopes and y intercepts are different,which in this problem they are, there is 1 solution. ========================== sub in (5,3) into the equations to see which comes out true.. D.) 2x - y = 7 2(5) - 3 = 7 10 - 3 = 7 7 = 7 (correct) 2x + 7y = 31 2(5) + 7(3) = 31 10 + 21 = 31 31 = 31 (correct) answer is D =========================== 8x - 3y = 1 -2x + 3y = 11 --------------add 6x = 12 x = 2 8x - 3y = 1 8(2) - 3y = 1 16 - 3y = 1 -3y = -16 + 1 -3y = -15 y = 5 solution is : (2,5) ==================== any questions ?

OpenStudy (love.kat123):

thank you